The electric field between the plates of charged parallel plate capacitor decreases from 12V/m to X V/m. If the number of atoms per unit volume is N x 1028 atoms m-3, the electronic polarisability of the material is
X = 6.00 N = 13.84
1
Expert's answer
2020-05-25T10:56:37-0400


Electronic polarisability , "\alpha_{e}"


"\alpha_{e} = \frac {\epsilon_{0}(\epsilon_{r}-1)} {N}"


we know that,


"X_{e} = \frac { \epsilon_{r}}{\epsilon_{0}}-1"


"6= \frac { 12}{\epsilon_{0}}-1"


"\epsilon_{0} = 1.7142"


Answer : "\alpha_{e} = \frac {\epsilon_{0}(\epsilon_{r}-1)} {N}"


"\alpha_{e} = \frac {1.7142(12-1)} {13.84\times 1028} = 1.3253\times10^{-3} F-m^{2}"